# Neural cyberattacks applied to the vision under realistic visual stimuli

### v1

@@ Abstract
Brain-Computer Interfaces (BCIs) are systems traditionally used in medicine and designed to interact with the brain to record or stimulate neurons. Despite their benefits, the literature has demonstrated that invasive BCIs focused on neurostimulation present vulnerabilities allowing attackers to gain control. In this context, neural cyberattacks emerged as threats able to disrupt spontaneous neural activity by performing neural overstimulation or inhibition. Previous work validated these attacks in small-scale simulations with a reduced number of neurons, lacking real-world complexity. Thus, this work tackles this limitation by analyzing the impact of two existing neural attacks, Neuronal Flooding (FLO) and Neuronal Jamming (JAM), on a complex neuronal topology of the primary visual cortex of mice consisting of approximately 230,000 neurons, tested on three realistic visual stimuli: flash effect, movie, and drifting gratings. Each attack was evaluated over three relevant events per stimulus, also testing the impact of attacking 25% and 50% of the neurons. The results, based on the number of spikes and shift percentages metrics, showed that the attacks caused the greatest impact on the movie, while dark and fixed events are the most robust. Although both attacks can significantly affect neural activity, JAM was generally more damaging, producing longer temporal delays, and had a larger prevalence. Finally, JAM did not require to alter many neurons to significantly affect neural activity, while the impact in FLO increased with the number of neurons attacked.
@@ 1 Introduction
Brain-Computer Interfaces (BCIs) are bidirectional systems that directly interact with the brain, allowing both neural data acquisition and neuromodulation, either by simulating or inhibiting neural activity. These technologies are classified as invasive, if it is necessary to implant electrodes directly into the brain, or non-invasive, if they use an electrode mesh over the scalp [1]. Non-invasive systems have been used thoroughly in the medical field for the diagnosis of neurological diseases, such as epilepsy [2]. However, these systems have gained special popularity in non-clinical scenarios, such as entertainment and video games, since they are accessible to the general public, present a low cost, and are portable. In contrast, invasive BCIs are useful for stimulating specific brain regions, being relevant for the treatment of neurodegenerative diseases, such as Parkinson’s disease [3]. Thus, companies such as Neuralink are focused on creating miniature implantable interfaces capable of offering broad neuronal coverage with low risk during surgery [4]. In this way, Neuralink currently investigates interfaces able to read and stimulate neurons individually to treat patients with neurological disorders, aiming to democratize the access to neurotechnology in the near future.
@@ 2 Related Work
This section first presents the state of the art oriented to broad aspects of cybersecurity on BCIs, showing potential attacks on confidentiality, integrity, and availability of BCIs, as well as users’ safety. After that, this section presents the works related to neural cyberattacks, highlighting the most relevant works and emphasizing important aspects of the solutions proposed, such as targeted brain areas and neuronal topologies used.
@@ 3.1 Solution design
This section presents the solution proposed to validate the impact of two neuronal cyberattacks, FLO and JAM, in a realistic neuronal topology. These attacks were selected because they are conceptually simple and have low complexity in implementation and execution. Additionally, they allow the stimulation and inhibition of neural activity, respectively, covering both attacking approaches. Figure 1 shows an overview of the proposed solution. First, the main component for building the topology is BMTK, which receives different stimuli as input offered by the authors of the realistic topology (1). The topology used in this work is proposed by Billeh et al. [27], who provide a complete reconstruction of a microcircuit from the primary visual cortex (V1) of mice. It comprises over 230,000 neurons with different populations and behaviors, including both excitatory and inhibitory neurons, modeling the characteristics of many morpho-electrical neuronal behaviors. Just like the topology from Arkhipov et al. [24] mentioned in Section 2, the topology used in this work provides two levels of granularity. The first level can simulate the behavior of the neurons in detail, and the electric and chemical processes that occur inside and between neurons (biophisiological approach), whereas the second level models the neuronal behavior according to the neuronal membrane potential in each instant (point-neurons). This point-neuron level comprises specific neurons called Gated Leaky Integrate and Fire (GLIF) neurons [28] that model neuronal behavior based on sets of differential equations of neuronal voltage over time. Both models must be created with the BMTK tool [25], and simulated accordingly to the approach, using NEURON [29] for the biophysically detailed model or NEST [26] for the point-neuron alternative. This research selected a point-neuron approach, which requires less computational resources and provides a sufficient approximation for the requirements of the investigation.
@@ 4.3 Attack parameters
The cyberattacks were configured taking into consideration different parameters to analyze the effectiveness of the attacks. These parameters have been already explored in the literature by López Bernal et al. [10, 9, 23] and López Madejska et al. [11]. This research considers as a parameter the number of neurons randomly attacked as specified in the literature, specifically selecting half (115,462) and a quarter (57,731) of the total number of neurons of the topology. Despite these parameters representing high percentages of the total number of neurons, these subsets are small compared to the whole of V1. Furthermore, each combination of attack parameters was executed ten times with the aim of analyzing if there existed a significant difference between distinct sets of randomly selected neurons. Additionally, the simulation duration is fixed at three seconds (3,000 ms) with a resolution of 0.25 ms (for evaluations per second), enough to assess the impact and propagation of the cyberattacks based on previous research [10].
Lastly, in FLO, this research establishes the voltage of the affected neurons to their individual threshold voltage (V_th). This specific voltage forces each neuron to generate sufficient activity to transmit an impulse to connected neurons. In JAM, the voltage of the neurons is forced to its minimum possible value (V_reset) to prevent the targeted neurons from having activity during the attack window. These parameters are selected after experiments performed by the literature [9] [11], where these works indicated that the ideal voltages to cause a greater effect are the values close to the threshold voltage in the case of FLO and the values near the minimum voltage in JAM. Furthermore, each combination of parameters was executed ten times to offer enough variability of randomly-selected neurons to obtain valid conclusions. Finally, Table 3 summarizes the parameters selected in this research for each cyberattack used.
@@ 5.1 Impact study over the flash effect
Starting with FLO, Figure 7 shows the behavior of the cyberattack over the flash effect, differentiating by the number of neurons targeted. The attack intervals are represented in the plots with a gray shade to visualize the duration of the cyberattack. The plots containing the attack over 25% of the neurons are represented on the left side (Figure 7(a) and Figure 7(c)), whereas the plots with 50% of the neurons under attack are on the right side (Figure 7(b) and Figure 7(d)). It is important to highlight that this distribution is maintained in all subsequent figures representing the results obtained. As observed in the plots, the attack instants selected (see Section 4.1) are distinguished by blue (ON-flash), pink (neutral event), and green (OFF-flash) lines in the case of the number of spikes metric, and bars for the shifts percentage metric. Besides, in the number of spikes metric, the lines are compared with the spontaneous behavior (orange line). Moreover, each line or bar contains the results for the ten executions performed per combination of attacking parameters, showing the average result per time interval.
@@
Figure 7(a) depicts the number of spikes analysis when the attack affects 25% of the total neurons. First, this figure shows that, in the ON-flash interval, there is an increase of 21,667 spikes (21,40%) compared to spontaneous behavior. On the contrary, attacking during the gray event, neural activity augments in 11,284 spikes (89.97%), whereas, in the OFF-flash, FLO has caused 10,651 spikes more (12,74%) over the spontaneous simulation. In contrast, Figure 7(b) represents the number of spikes when 50% of neurons are attacked. In the first event, the attack causes an increase of 52,742 spikes (52,08%), although the augmentation in the neutral event is 28,815 spikes (229,7%). Finally, in the OFF-flash, there have been 27,444 additional spikes (32,83%) over the spontaneous signaling. As can be seen, the gray event is the most impacted as neurons spontaneously exhibit a lower response than during other events; thus, any malicious overstimulation will cause a notorious variation. Nevertheless, it is important to highlight that the attack impact can vary depending on the stimulus presented since neurons naturally behave in different ways according to visual inputs. Moreover, after attacking each event, the impact regarding the number of spikes metric suffers a propagation that lasts around two or three intervals (200-300 ms) until the neurons return to normal behavior. The recovery of the spontaneous state occurs in all experiments as the topology is composed of realistic individual neurons, which tend to synchronize again with their normal activity based on the visual stimulus.
@@
Regarding the analysis of shifts percentages, Figure 7(c) represents the results when the attacks affect 25% of the neurons. First of all, attacking during the ON-flash causes a 61% shift increase. However, attacking during the gray event forces 94% of the spikes to be delayed, while the attack during the OFF-flash generates 16% of shifts. In contrast, the percentage of shifts when 50% of neurons are affected is represented on Figure 7(d). Attacking during the ON-flash causes a variation of 71%, whereas the neutral event and OFF-flash obtain 98% and 29% shifts, respectively. As observed, the gray event obtained the higher percentages as in the last metric. Nevertheless, there is a substantial increase in all intervals under attack, followed by fluctuations maintained until the end of the simulation (around 40%-60%). All these results indicate that, although neurons seem to return to the spontaneous activity according to the number of spikes metric, the attack causes a delay of the activations in time that gets propagated over the whole simulation, highlighted by the maintained effect observed by this metric.
@@
Figure 8 represents the impact of JAM when inhibiting different time windows selected in Section 4.1. Starting with Figure 8(a), where the 25% of the neurons are inhibited, the ON-flash presents a mean decrease of 95,696 spikes (-94.50%). In contrast, attacking during the neutral interval reaches a decrement of 11,910 spikes (-94.96%), while the impact for the OFF-flash decreases the number of spikes by 77,413 spikes (-92.60%) compared to spontaneous behavior. Next, Figure 8(b) shows that the attack produces a decrease of 99,461 spikes (-98.22%) in the white event, while in the gray event the number of spikes decreases by 12,290 spikes (-97.99%). Finally, the OFF-flash there have been 81,794 spikes less (-97.85%). Once JAM has finished, the inhibited neurons generate a high peak of activations, caused by the liberation of the neurons after the attack, which tend to return to their normal behavior in a coordinated way. The results highlight the similarity between attacking both number of neurons, implying that it is not necessary to inhibit a large set of neurons to produce a significant impact. Depending on the event the attack affects, the propagation of the impact differs. Attacking the ON-flash and gray events takes around 11 intervals (1,100 ms) to converge. In contrast, in the OFF-flash, the neurons take seven intervals (700ms) to return to normal activity.
@@
In the percentage of shifts metric, represented by Figure 8(c) for attacking 25% of the neurons, the attack interval of the ON-flash reaches 33% of shifts, the gray event 16%, and the OFF-flash 48% of shifts. Oppositely, Figure 8(d) depicts that the white event obtains a percentage of 12%. In contrast, the neutral event reaches a percentage of 4%, while the black event obtains 27% of shifts. As can be observed, the reduction in the number of spikes directly affects the percentage of shifts since having a very low number of spikes per interval will restrict the variability of the spikes over time. Moreover, the high peak produced after the cyberattack, already explained in the previous metric, is also represented in this metric after the gray shades. More particularly, the intervals after the attacks start to fluctuate until the percentages stabilize between 50%-70%.
(a) Impact on the number of spikes over 25% of the neurons.
(b) Number of spikes analysis when the attack affects 50% of the neurons.
(c) Shifts percentages caused by the attack when targeting 25% of the neurons.
(d) Impact on the percentage of shifts over 50% of the neurons.
@@
(a) Number of spikes when JAM inhibits 25% of the neurons.
(b) Number of spikes when targeting 50% of the neurons.
(c) Shifts analysis when affecting 50% of the neurons.
(d) Shifts percentages of the impact over 50% of the neurons.

### v2

@@ Abstract
Brain-Computer Interfaces (BCIs) are systems traditionally used in medicine and designed to interact with the brain to record or stimulate neurons. Despite their benefits, the literature has demonstrated that invasive BCIs focused on neurostimulation present vulnerabilities allowing attackers to gain control. In this context, neural cyberattacks emerged as threats able to disrupt spontaneous neural activity by performing neural overstimulation or inhibition. Previous work validated these attacks in small-scale simulations with a reduced number of neurons, lacking real-world complexity. Thus, this work tackles this limitation by analyzing the impact of two existing neural attacks, Neuronal Flooding (FLO) and Neuronal Jamming (JAM), on a complex neuronal topology of the primary visual cortex of mice consisting of approximately 230,000 neurons, tested on three realistic visual stimuli: flash effect, movie, and drifting gratings. Each attack was evaluated over three relevant events per stimulus, also testing the impact of attacking 25% and 50% of the neurons. The results, based on the number of spikes and shift percentages metrics, showed that the attacks caused the greatest impact on the movie, while dark and static events exhibit highest resilience. Although both attacks can significantly affect neural activity, JAM was generally more damaging, producing longer temporal delays, and had a larger prevalence. Finally, JAM did not require to alter many neurons to significantly affect neural activity, while the impact in FLO increased with the number of neurons attacked.
@@
{highlights}
Evaluation of two neural cyberattacks on a realistic simulation of the visual cortex.
Use of a rich topology with 230,000 neurons from six cortical layers.
Attacks tested on different events within three visual stimuli used as visual input.
Evaluation of both threats using two metrics: number of spikes and shifts percentage.
Both attacks can disrupt neural activity, having effect for hundreds of milliseconds.
@@ 1 Introduction
Brain-Computer Interfaces BCIs are bidirectional systems that directly interact with the brain, allowing both neural data acquisition and neuromodulation, either by simulating or inhibiting neural activity. These technologies are classified as invasive, if it is necessary to implant electrodes directly into the brain, or non-invasive, if they use an electrode mesh over the scalp [1]. Non-invasive systems have been used thoroughly in the medical field for the diagnosis of neurological diseases, such as epilepsy [2]. However, these systems have gained special popularity in non-clinical scenarios, such as entertainment and video games, since they are accessible to the general public, present a low cost, and are portable. In contrast, invasive BCIs are useful for stimulating specific brain regions, being relevant for the treatment of neurodegenerative diseases, such as Parkinson’s disease [3]. Thus, companies such as Neuralink are focused on creating miniature implantable interfaces capable of offering broad neuronal coverage with low risk during surgery [4]. In this way, Neuralink currently investigates interfaces able to read and stimulate neurons individually to treat patients with neurological disorders, aiming to democratize the access to neurotechnology in the near future.
@@ 2 Related Work
This section first presents the neural models studied from the literature and their characteristics. Next, this section descibes the state of the art oriented to broad aspects of cybersecurity on BCIs, showing potential attacks on confidentiality, integrity, and availability of BCIs, as well as users’ safety. After that, this section presents the works related to neural cyberattacks, highlighting the most relevant works and emphasizing important aspects of the solutions proposed, such as targeted brain areas and neuronal topologies used.
2.1 Reconstruction of realistic neural models.
There are numerous models focused on different aspects of the brain. These models are realistic but on a small scale. Moreover, most of them are not sufficiently complex or complete to study the dynamics of neural cyberattacks against neural activity. Therefore, this work reviews models that realistically simulate and replicate neural behavior, presenting sufficient details. Nevertheless, for the sake of simplicity, this section highlights the most relevant models for this research. The selection criteria focus on models providing a complete reconstruction of a specific brain microcircuit, offering a high degree of biological realism, and if they are publicly available for research purposes.
[…]
Aussel et al. [14] created a computational model of the healthy hippocampus (CA1, CA3, and dentate gyrus) as well as the entorhinal cortex , which presents characteristic rhythms of wakefulness and slow-wave sleep. This network comprises 33,200 neurons, including 31,000 excitatory neurons and 2,200 inhibitory interneurons, distributed across the following components: CA1 (10,000 excitatory neurons; 1,000 interneurons), CA3 (1,000; 100), dentate gyrus (10,000; 100), and entorhinal cortex (10,000; 1,000). The authors utilized point-neurons with realistic dynamics, and conductance-based Hodgkin-Huxley neurons. The results showed that this model can reproduce two types of oscillations observed in humans: sharp-wave ripples during deep sleep and theta-nested gamma oscillations during wakefulness. Later, the same authors extended the previous model of a healthy hippocampus, including four typical pathological modifications of the hippocampus seen in medial temporal lobe epilepsy [15]. This extension aimed to investigate how these structural and functional abnormalities, both at the neuronal and network levels, contribute to the generation of seizure-like theta activity under different brain states (wakefulness and slow-wave sleep).
Finally, Romani et al. [16] presented a full-scale model of the rat hippocampus, particularly from CA1, comprising 456,000 neurons derived from 43 morphological reconstructions of neurons belonging to 12 distinct morphological types. The authors focused on the reconstruction and validation of the model, followed by the simulation and analysis of activity in this reconstructed topology. Additionally, the input used in this work is, mainly, the Schaffer collaterals from CA3 pyramidal cells, the most prominent afferent input to the CA1 and the most studied pathway in the hippocampus. They concluded that this model is one of the most detailed and comprehensive of this region to date, capable of reproducing realistic dynamics. However, it still presents limitations, as the model has incomplete structures due to a lack of cell classes, data, and standardized validation.
@@ 3.1 Solution design
This section presents the solution proposed to validate the impact of two neuronal cyberattacks, FLO and JAM, in a realistic neuronal topology. These attacks were selected because they are conceptually simple and have low complexity in implementation and execution, unlike other neural cyberattacks that require a more challenging implementation. Additionally, they allow the stimulation and inhibition of neural activity, respectively, covering both attacking approaches. They also differ in terms of duration and temporal resolution: one produces an instantaneous effect, while the other operates over a defined time window. This contrast allows the evaluation of different attack propagation mechanisms. Figure 1 shows an overview of the proposed solution. First, the main component for building the topology is BMTK, which receives different stimuli as input offered by the authors of the realistic topology (1). The topology used in this work is proposed by Billeh et al. [13], who provide a complete reconstruction of a microcircuit from the primary visual cortex (V1) of mice. It comprises over 230,000 neurons with different populations and behaviors, including both excitatory and inhibitory neurons, modeling the characteristics of many morpho-electrical neuronal behaviors. This topology was chosen due to its high biological realism, as it not only integrates detailed morpho-electrical properties and synaptic connectivity, but also reconstructs the full six-layer architecture of mice V1. Compared to other available models, it offers the most comprehensive representation of visual processing, capturing both structural and functional features of the visual cortical microcircuit with a high degree of precision. Just like the topology from Arkhipov et al. [12] mentioned in Section 2, the topology used in this work provides two levels of granularity. The first level can simulate the behavior of the neurons in detail, and the electric and chemical processes that occur inside and between neurons (biophisiological approach), whereas the second level models the neuronal behavior according to the neuronal membrane potential in each instant (point-neurons). This point-neuron level comprises specific neurons called Gated Leaky Integrate and Fire (GLIF) neurons [31] that model neuronal behavior based on sets of differential equations of neuronal voltage over time. Both models must be created with the BMTK tool [29], and simulated accordingly to the approach, using NEURON [32] for the biophysically detailed model or NEST [30] for the point-neuron alternative. This research selected a point-neuron approach, which requires less computational resources and provides a sufficient approximation for the requirements of the investigation. This selection was based on the high computational cost associated with biophysically detailed models. In contrast, the point-neuron approach provides a sufficiently realistic representation of neural behavior without requiring detailed structural information about neurons. Additionally, it enables faster simulations, making it an efficient solution. Furthermore, analyzing the impact of cyberattacks does not require a highly detailed biophysical model, as the point-neuron approach proves sufficient for observing and evaluating their effects.
@@ 4.3 Attack parameters
The cyberattacks were configured taking into consideration different parameters to analyze the effectiveness of the attacks. These parameters have been already explored in the literature by López Bernal et al. [10, 9, 28] and López Madejska et al. [11]. One of the key aspects of the configuration is that the cyberattacks are applied across all six layers of the V1 model. This decision ensures that the impact of the attacks is distributed throughout the entire cortical structure, rather than being limited to a specific region or subset of neurons. By targeting the entire visual cortex, the evaluation provides a more comprehensive understanding of the global effects induced by each attack. This research considers as a parameter the number of neurons randomly attacked as specified in the literature, specifically selecting half (115,462) and a quarter (57,731) of the total number of neurons of the topology. Despite these parameters representing high percentages of the total number of neurons, these subsets are small compared to the whole of V1. Furthermore, each combination of attack parameters was executed ten times with the aim of analyzing if there existed a significant difference between distinct sets of randomly selected neurons. Additionally, the simulation duration is fixed at three seconds (3,000 ms) with a resolution of 0.25 ms (for evaluations per second), enough to assess the impact and propagation of the cyberattacks based on previous research [10].
Lastly, in FLO, this research establishes the voltage of the affected neurons to their individual threshold voltage (V_th). This specific voltage forces each neuron to generate a spike, thereby ensuring the effectiveness of the attack. This voltage provokes the neuron to transmit an impulse to those neurons with which it has synapses. In JAM, the voltage of the neurons is forced to its minimum possible value, corresponding to the reset potential (V_reset) to prevent the targeted neurons from having activity during the attack window. This abrupt reduction effectively maintains the neurons below their firing threshold, thereby suppressing their ability to generate an action potential. These parameters are selected after experiments performed by the literature [9] [11], where these works indicated that the ideal voltages to cause a greater effect are the values close to the threshold voltage in the case of FLO and the values near the minimum voltage in JAM. Furthermore, each combination of parameters was executed ten times to offer enough variability of randomly-selected neurons to obtain valid conclusions. Finally, Table 3 summarizes the parameters selected in this research for each cyberattack used.
@@ 5.1 Impact study over the flash effect
Starting with FLO, Figure 7 shows the behavior of the cyberattack over the flash effect, differentiating by the number of neurons targeted. The attack intervals are represented in the plots with a gray shade and red numbers to visualize the duration of the cyberattack. The plots containing the attack over 25% of the neurons are represented on the left side , whereas the plots with 50% of the neurons under attack are on the right side . It is important to highlight that this distribution is maintained in all subsequent figures representing the results obtained. As observed in the plots, the attack instants selected (see Section 4.1) are distinguished by blue (ON-flash), pink (neutral event), and green (OFF-flash) lines in the case of the number of spikes metric, and bars for the shifts percentage metric. Besides, in the number of spikes metric, the lines are compared with the spontaneous behavior (orange line). Moreover, each line or bar contains the results for the ten executions performed per combination of attacking parameters, showing the average result per time interval.
@@
Figure 7(a), the left plot, depicts the number of spikes analysis when the attack affects 25% of the total neurons. First, this figure shows that, in the ON-flash interval, there is an increase of 21,667 spikes (21.40%) compared to spontaneous behavior. On the contrary, attacking during the gray event, neural activity augments in 11,284 spikes (89.97%), whereas, in the OFF-flash, FLO has caused 10,651 more spikes (12.74%) over the spontaneous simulation. In contrast, the right figure represents the number of spikes when 50% of neurons are attacked. In the first event, the attack causes an increase of 52,742 spikes (52.08%), although the augmentation in the neutral event is 28,815 spikes (229.7%). Finally, in the OFF-flash, there have been 27,444 additional spikes (32.83%) over the spontaneous signaling. As can be seen, the gray event is the most impacted as neurons spontaneously exhibit a lower response than during other events; thus, any malicious overstimulation will cause a notorious variation. Nevertheless, it is important to highlight that the attack impact can vary depending on the stimulus presented since neurons naturally behave in different ways according to visual inputs. Moreover, after attacking each event, the impact regarding the number of spikes metric suffers a propagation that lasts around two or three intervals (200-300 ms) until the neurons return to normal behavior. The recovery of the spontaneous state occurs in all experiments as the topology is composed of realistic individual neurons, which tend to synchronize again with their normal activity based on the visual stimulus.
@@
Regarding the analysis of shifts percentages, the left side of Figure 7(b) represents the results when the attacks affect 25% of the neurons. First of all, attacking during the ON-flash causes a 61% shift increase. However, attacking during the gray event forces 94% of the spikes to be delayed, while the attack during the OFF-flash generates 16% of shifts. In contrast, the percentage of shifts when 50% of neurons are affected is represented on the right side of Figure 7(b). Attacking during the ON-flash causes a variation of 71%, whereas the neutral event and OFF-flash obtain 98% and 29% shifts, respectively. As observed, the gray event obtained the higher percentages as in the last metric. Nevertheless, there is a substantial increase in all intervals under attack, followed by fluctuations maintained until the end of the simulation (around 40%-60%). All these results indicate that, although neurons seem to return to the spontaneous activity according to the number of spikes metric, the attack causes a delay of the activations in time that gets propagated over the whole simulation, highlighted by the maintained effect observed by this metric.
@@
Figure 8 represents the impact of JAM when inhibiting different time windows selected in Section 4.1. Starting with the left plot on Figure 8(a), where the 25% of the neurons are inhibited, the ON-flash presents a mean decrease of 95,696 spikes (-94.50%). In contrast, attacking during the neutral interval reaches a decrement of 11,910 spikes (-94.96%), while the impact for the OFF-flash decreases the number of spikes by 77,413 spikes (-92.60%) compared to spontaneous behavior. Next, the right figure shows that the attack produces a decrease of 99,461 spikes (-98.22%) in the white event, while in the gray event the number of spikes decreases by 12,290 spikes (-97.99%). Finally, the OFF-flash there have been 81,794 spikes less (-97.85%). Once JAM has finished, the inhibited neurons generate a high peak of activations, caused by the liberation of the neurons after the attack, which tend to return to their normal behavior in a coordinated way. The results highlight the similarity between attacking both number of neurons, implying that it is not necessary to inhibit a large set of neurons to produce a significant impact. Depending on the event the attack affects, the propagation of the impact differs. Attacking the ON-flash and gray events takes around 11 intervals (1,100 ms) to converge. In contrast, in the OFF-flash, the neurons take seven intervals (700ms) to return to normal activity.
@@
In the percentage of shifts metric, represented by Figure 8(b), the left plot for attacking 25% of the neurons, the attack interval of the ON-flash reaches 33% of shifts, the gray event 16%, and the OFF-flash 48% of shifts. Oppositely, the right side of Figure 8(b) depicts that the white event obtains a percentage of 12%. In contrast, the neutral event reaches a percentage of 4%, while the black event obtains 27% of shifts. As can be observed, the reduction in the number of spikes directly affects the percentage of shifts since having a very low number of spikes per interval will restrict the variability of the spikes over time. Moreover, the high peak produced after the cyberattack, already explained in the previous metric, is also represented in this metric after the gray shades. More particularly, the intervals after the attacks start to fluctuate until the percentages stabilize between 50%-70%.
(a) Impact on the number of spikes over 25% (left figure) and 50% (right figure) of the neurons. The red numbers indicated the three events under attack individually, specified in Section 4.1.
(b) Impact on the percentage of shifts over 25% (left figure) and 50% (right figure) of the neurons. Events attacked individually are indicated with red numbers.
@@
(a) Number of spikes when JAM inhibits 25% (left figure) and 50% (right figure) of the neurons, on the three events indicated with gray shades and red numbers.
(b) Shifts percentages of the impact over 25% (left figure) and 50% (right figure) of the neurons. Red numbers and gray shades highlighted the events under attack.